OpenStudy (istim):

he following definite integral can be evaluated by subtracting F(B) - F(A), where F(B) and F(A) are found from substituting the limits of integration. \[\int_{0}^{4} \frac{4000 x +400 }{(5 x^2 +1 x +4)^5}dx\] After substitution, the upper limit of integration (B) is : and the lower limit of integration (A) is : After integrating, F(B) = F(A) =

OpenStudy (anonymous):

You have a polynomial raised to a power in the denominator of the integrand so choosing that polynomial as your substitution would be a nice place to start.

OpenStudy (istim):

I know B is 88 and A is 4. I need help with the last 2.

OpenStudy (anonymous):

What is the anti-derivative of the integrand after substitution? Find that and then use the new limits you got (A and B) to find F(A) and F(B).

OpenStudy (callisto):

\[\int_{0}^{4} \frac{4000 x +400 }{(5 x^2 +1 x +4)^5}dx=\int_{0}^{4} \frac{400(10 x +1) }{(5 x^2 +1 x +4)^5}dx=\int_{0}^{4} \frac{400}{(5 x^2 +1 x +4)^5}(10 x +1) dx\] Let u = \(5x^2+x+4\) , du = (10x+1)dx. Then the integral becomes \[\int_{a'}^{b'} \frac{400 }{(u)^5}du\] ,where a' and b' are the new limits.

OpenStudy (istim):

u=5x^2+1x+4. Du=10x+1 400du=4000x+400

OpenStudy (callisto):

First, integrate, then substitute.

OpenStudy (istim):

400du/u^5?

OpenStudy (anonymous):

Yes, integrate that and use the limits of integration you found earlier.

OpenStudy (istim):

400du/(1.6u^6)?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Power_rule

OpenStudy (istim):

differentiate, no integrate? or am i just to tired

OpenStudy (istim):

hello

OpenStudy (istim):

@JohnathanHocker or@Callisto ?

OpenStudy (nincompoop):

have you learned u-substitution yet?

OpenStudy (istim):

yeah, thats what i thought it was

OpenStudy (nincompoop):

what is the goal when doing u-sub?

OpenStudy (istim):

no idea, sorry.

OpenStudy (istim):

I just did it.

OpenStudy (nincompoop):

k it is du = 10x+1 dx

OpenStudy (istim):

Yes, I know thus far.

OpenStudy (nincompoop):

and when doing u-substitution you have to check for a and b values because the range might not be from 0 - 4 anymore

OpenStudy (istim):

41 and 1?

OpenStudy (nincompoop):

did you solve or guess it?

OpenStudy (istim):

subbed into the equations derived

OpenStudy (nincompoop):

when x = 0 u = ? when x = 4 u = ? that is how you do it then use the new values also, notice now that you have two 10x+1? the one callisto did will end up being as the "dx" when you finally use the u notation

OpenStudy (istim):

so did my values work?

OpenStudy (nincompoop):

dunno

OpenStudy (istim):

oh.

OpenStudy (istim):

I'm not subbing into the derived equation right?

OpenStudy (istim):

@nincompoop ?

OpenStudy (nincompoop):

you are use the expression you're going to change into u then compare it to the u

OpenStudy (istim):

whereas u was 5x^2+x+4?

OpenStudy (istim):

@nincompoop @Callisto @JonathanHocker any ideas?

OpenStudy (anonymous):

@IsTim what you don't get?

OpenStudy (istim):

To be honest, I feel like I don't understand anything. Lots of people online and offline have explained this to me, but I keep getting lost.

OpenStudy (istim):

anoyne?

OpenStudy (anonymous):

so you don't get what @callisto did?

OpenStudy (istim):

I managed to get as far as where he stated, but I don't know what to do after (Adn I still don't understand what I did up to that point)

OpenStudy (anonymous):

F(x)=\[\int_{0}^{4} \frac{4000 x +400 }{(5 x^2 +1 x +4)^5}dx\] \[= 400\int\limits_{0}^{4} \frac{ 10x+1 }{ (5x^2 +x+4 )^5}dx\]

OpenStudy (istim):

So we forget out 400?

ganeshie8 (ganeshie8):

^^ we can do that because :- \(\large \int cf(x) dx = c \int f(x) dx\)

ganeshie8 (ganeshie8):

u can pull out 'constant' out of integral

OpenStudy (anonymous):

yes only focus on terms inside in integral sign

ganeshie8 (ganeshie8):

Also, just for completeness :- \(\large \int c dx = c \int 1 dx = cx + c_1\)

ganeshie8 (ganeshie8):

u need to knw these basic properties before diving into doing complicated indefinite integrals

OpenStudy (anonymous):

for make easy this integration we let 5x^2+x+4=u and differentiate it as du/dx = 10x +1 which can be written as du=(10x+1)dx

OpenStudy (istim):

I know what you're talking about; I've achieved the same results, I jsut don't know ti use this to figure out constant of integration

ganeshie8 (ganeshie8):

actually we're doing a "definite integral", the constant gets subtracted away when u do : F(B) - F(A) So, we dont bother about constant of integration in "definite integrals"

OpenStudy (istim):

Oh no, I forgot to include that in this question. They're looking for that.

OpenStudy (istim):

So sorry. I reviewed the question and realized I was missing something.

OpenStudy (istim):

I was wonderin why everyone was telling me what I already know, so I looked through and realized I forgot to ask for constant of integration

OpenStudy (istim):

OH MY GOD SORRY! LAck of sleep. IT isn't asking for that. I'm doing 2 questions at the same.time. Ignore constant of something.

OpenStudy (istim):

It's: After integrating, F(B) = ? F(A)=?

OpenStudy (istim):

That's what I can't figure out.

OpenStudy (callisto):

After integrating, you can put the value of the upper limit to get F(B), and the value of the lower limit to get F(A)

OpenStudy (callisto):

Perhaps you can forget about the definite integral, and do the indefinite integral, shown as below, first: \[400\int\limits\frac{ 10x+1 }{ (5x^2 +x+4 )^5}dx\]

OpenStudy (istim):

not gettin the right answers though.

OpenStudy (istim):

F(B)=-1.66 and F(A)=-0.39

OpenStudy (istim):

Both in terms of charge and value.

OpenStudy (istim):

Hello?

OpenStudy (istim):

I got 0.0226 and 0.465

OpenStudy (callisto):

Would you mind showing us what you have got for the integral BEFORE evaluating it?

OpenStudy (callisto):

That was mechanics...

OpenStudy (istim):

sorry i don't understand?

OpenStudy (istim):

wait darnit wrong picture

OpenStudy (istim):

goddarnit i need sleep

OpenStudy (callisto):

I guess you have not integrated it yet. You just substituted the values in it.

OpenStudy (callisto):

1) You need to integrate\[400\int\limits\frac{ 10x+1 }{ (5x^2 +x+4 )^5}dx\]as usual, the final answer should be in terms of x. 2) Substitute the values (the upper limit and lower limit) back to the result you get. You can drop the constant you got when doing the substitution.

OpenStudy (istim):

Does that work?

OpenStudy (callisto):

Of course, it does not.

OpenStudy (callisto):

As mentioned above, you need to use u-substitution.

OpenStudy (istim):

400lnu^5?

OpenStudy (istim):

c is -2768?

OpenStudy (callisto):

Not ln(u^5). Use power rule when you have \[\int \frac{1}{x^a}dx\]for a\(\ne\)1.

OpenStudy (istim):

Sorry, you've confused me a bit more.

OpenStudy (istim):

argh everyone always leaves

OpenStudy (istim):

@Callisto @ganeshie8 @gyanu @nincompoop @JonathanHorker any ideas what that means/how to use it?

OpenStudy (callisto):

\[\int x^adx =\frac{1}{a+1}x^{a+1}+C, \text{where } a\ne1 \]

OpenStudy (istim):

a is dx?

OpenStudy (callisto):

a is a constant, the power of x.